4. Based on the method presented in the lesson DIFFEQ 017, provide the derivation of the Laplace transform L{sin at.

Solution steps plz
Answer #1

o Sin lat)e-st dt I- - Using Definition 2(f(+) J f (4) e-st.dt d(sincet)= si J Sin/at) e-87 at = sin(at)e-st facosat = sin loÓ I =fsin (at) e-st det Sinlat)est 4 - a Coslat e-st lt az a la ood(sin ct) = 5 Ginlat)e-st dt = Sin (at)e-st. a alos lat)e-8